Manuel is 25 years old and his brother is 19. How many years ago was Manuel four times as old as David. I need help with how to set up the equation.

Answer by stanbon(75887)

Manuel is 25 years old and his brother is 19. How many years ago was Manuel four times as old as David.
-----
Assume the answer is "x" years.
----------
Equation:
25-x = 4(19-x)
------
25-x = 76 - 4x
3x = 51
x = 17 years
==============
Cheers,
Stan H.
---------------